When 14g of an impure sample of iron was dissolved in 300cm^3 of HCL solution the mass of hydrogen produced was 0.4g calculate the percentage purity of the iron. The mass of FeCl2 formed and the concentration of HCL used.


1
Expert's answer
2021-03-16T07:49:01-0400

m(sample) = 14 g

V(HCl) = 300 ml = 0.3 L

m(H2) = 0.4 g


x g 0.4 g

Fe + 2HCl = FeCl2 + H2

55,847 g 2 g

x/55,847 = 0.4/2

x = 11.17 g

w(Fe) = 11.17 g / 14 g * 100% = 79.8%


y mol 0.4 g

Fe + 2HCl = FeCl2 + H2

2 mol 2 g

y/2 = 0.4/2

y = 0.4 mol

C(HCl) = n(HCl)/V(HCl) = 0.4 mol / 0.3 L = 1.33 mol/L


z g 0.4 g

Fe + 2HCl = FeCl2 + H2

126.75 g 2 g

z/126.75 = 0.4/2

z = 25.34 g FeCl2